1.    Calculate the pH of a buffer solution consisting of 0.50 M HC2H3O2 and 0.50 M NaC2H3O2 (buffer solution A). Ka of HC2H3O2 = 1.75 x 10-5.

a.     Calculate the new pH of the buffer solution after adding 0.020 mol HCl to 1.0 L of buffer solution A (5 pts).

b.    Calculate the new pH of the buffer solution after adding 0.020 mol NaOH to 1.0 L of buffer solution A (5 pts).


1
Expert's answer
2022-05-18T13:34:03-0400

pH = pKa + log ([salt]/[acid])

pKa = - logKa = 4.76

pH = 4.76 + log(0.5/0.5) = 4.76


a) pH2 = pKa + log ((0.5-0.02)/(0.5+0.02))= 4.76 - 0.03 = 4.73


b) pH3 = pKa + log(0.5+0.02)/(0.5-0.02) = 4.76 + 0.03 = 4.79